### v3

# Autocomp: LLM-Driven Code Optimization for Tensor Accelerators

Abstract
Hardware accelerators, especially those designed for tensor processing, have become ubiquitous in today’s computing landscape. However, even with significant efforts in building compilers, programming these tensor accelerators remains challenging, leaving much of their potential underutilized. Recently, large language models (LLMs), trained on large amounts of code, have shown significant promise in code generation and optimization tasks, but generating low-resource languages like specialized tensor accelerator code still poses a significant challenge. We tackle this challenge with Autocomp, an approach that empowers accelerator programmers to leverage domain knowledge and hardware feedback to optimize code via an automated LLM-driven search. We accomplish this by: 1) formulating each optimization pass as a structured two-phase prompt, divided into planning and code generation phases, 2) inserting domain knowledge during planning via a concise and adaptable optimization menu, and 3) integrating correctness and performance metrics from hardware as feedback at each search iteration. Across three categories of representative workloads and two different accelerators, we demonstrate that Autocomp-optimized code runs 5.6 (GEMM) and 2.7 (convolution) faster than the vendor-provided library, and outperforms expert-level hand-tuned code by 1.4 (GEMM), 1.1 (convolution), and 1.3 (fine-grained linear algebra). Additionally, we demonstrate that optimization schedules generated from Autocomp can be reused across similar tensor operations, improving speedups by up to 24% under a fixed sample budget.
1 Introduction
Hardware accelerators [43, 29] have become a critical driving force for the recent breakthroughs [32, 18, 19, 59, 44] in machine learning. They provide hundred-fold improvements in performance and energy efficiency in running deep neural networks (DNNs), and this has led to an increasing number of accelerators for tensor processing in recent years [29, 46, 42, 34, 27, 13, 6, 8]. However, extracting that performance requires writing high-performance accelerator code, which is time consuming and requires a deep understanding of the underlying hardware.
To address this challenge, various compilers and domain-specific languages (DSLs) have appeared. For deep learning applications, compilers such as XLA, TVM, and Triton generate high-performance code, but they only support a few hardware backends, particularly CPUs and GPUs [45, 7]. Unfortunately, adapting compilers and DSLs to new hardware platforms with vendor-specific instruction set architectures (ISAs) and implementation-specific dataflow patterns requires significant engineering effort. In fact, software alone comprises 40-50% of the development cost for new hardware [58, 54, 23], even before considering the effort needed for end users to write and debug software for a newly developed chip. Prior work in DSLs like Halide and Exo [50, 26] targets accelerators by providing primitives that make it easier to express tensor computation, but the onus of optimizing code written in such DSLs still lies on the accelerator programmer.
Even once a compiler exists, generating performant code runs into the classical “scheduling” problem, i.e., deciding which optimizations to apply and in what order. For general-purpose backends (CPUs and GPUs), these passes have been iteratively developed and refined over many years through a combination of experts and auto-tuning frameworks. Recent work has gone further in exploring data-driven approaches such as supervised learning [66] and reinforcement learning [9], and even LLM training [10] to tackle the combinatorial explosion of pass sequences. While these data-driven approaches have shown promise, they depend on vast amounts of performance data to train, which is painfully scarce for domain-specific hardware accelerators.
In this paper, we present Autocomp, which solves the problems with prior approaches with an iterative LLM-guided search framework to optimize accelerator code. Unlike previous compilers, Autocomp can adapt to new hardware platforms and ISAs by simply changing prompts. Unlike existing tensor DSLs, Autocomp automatically generates optimized code without manual tuning. And unlike data-driven approaches targeting CPUs and GPUs, Autocomp requires no model training, instead leveraging LLM in-context reasoning and pretrained knowledge of common optimizations.
In each iteration, Autocomp first plans by choosing an optimization from a predefined menu, i.e., a list of common hardware accelerator optimizations like tiling and unrolling, then applies the optimization to generate optimized DSL code. The generated candidates are validated for correctness and benchmarked on the accelerator to collect performance metrics, providing feedback for the next iteration of search. By encoding DSL syntax, optimization rules, and performance feedback concisely in a prompt, Autocomp can guide the LLM to generate optimized code to run on the target accelerator.
In our evaluation, we apply Autocomp to two representative low-resource accelerators and generate code that runs 5.6 (GEMM) and 2.7 (convolution) faster than the vendor-provided library. Furthermore, it outperforms expert-level hand-tuned code by 1.4 (GEMM), 1.1 (convolution), and 1.3 (fine-grained linear algebra), surpassing the prior best known implementations with less human effort. Moreover, we show that Autocomp’s schedules can be reused as guidance when scheduling similar tensor operations, alleviating the cost of scheduling new code and delivering up to 24% greater speedups under a fixed sample budget.
In summary, we make the following contributions in this paper:
1.
We present Autocomp, the first LLM-driven code optimization approach for low-resource tensor accelerator code generation.
2.
Autocomp incorporates domain knowledge, hardware feedback on correctness and performance, and novel strategies for response diversity to automatically generate performant code.
3.
Our generated implementations significantly outperform hand-optimized code written by experts across a wide range of workloads and across different tensor accelerators.
4.
We illustrate that schedules generated by Autocomp can be reused to optimize similar tensor operations, reducing search cost and demonstrating the a posteriori usefulness of Autocomp-generated schedules beyond pure performance.
2 Background
2.1 Code Optimization for Tensor Accelerators
Programming tensor accelerators differs greatly from programming general-purpose CPUs. Tensor accelerators, depicted in Fig. 2, generally focus on the efficient execution of fixed-size (e.g., 1616) matrix multiplication instructions, as shown in Fig. 2. Rather than trying to reduce the number or type of these instructions, which is often fixed, software optimization focuses on other aspects, such as:
Minimizing data movement between main memory and smaller accelerator-local memories (in Fig. 2, the scratchpad and accumulator).
Setting configuration state for computation and data movement.
Scheduling or reordering operations to maximally overlap computation and data movement.
Code transformations that enable these optimizations range from low-level changes like arithmetic simplification or instruction selection, to higher-level changes like loop tiling, hoisting (Fig. 4), or software pipelining (Fig. 4). These high-level changes, while improving performance, require loop nests, pointers, and indices to be modified in multiple locations, making them challenging to implement, especially in a low-resource DSL.
Prior work has explored some of this optimization space. For example, performance models like Timeloop [48], MAESTRO [33], and TeAAL [39] use high-level hardware architectural models and software abstractions to represent tensor accelerators and their workloads. Much recent work has sought to automatically explore this space, using methods such as machine learning [21, 25, 53, 20], linear programming [24], black-box optimization [30, 65, 52], and reinforcement learning [64]. While these abstractions capture some aspects of tensor accelerator code optimization, in particular the amount of data movement, they neglect other implementation-specific and instruction-level optimizations. In this work, LLM code generation allows us to directly rewrite accelerator code, expanding the search to include all potential axes of optimization.
2.2 LLM-Based Code Optimization
LLMs have been used in various code-related tasks [5, 41, 2]. For code optimization, several works have utilized evolutionary approaches that take advantage of LLMs’ ability to apply novel mutations to code [36, 4, 38, 37, 16, 51]. Others have used methods like in-context learning and retrieval-augmented generation (RAG) to transfer knowledge from a library of previously known optimizations to the current problem [61, 1]. Still others have collected large datasets of performance examples to fine-tune models [55], construct simplified program abstractions for faster optimization [62], or utilize compiler and runtime feedback to iteratively improve the generated code [49, 11, 57, 47].
Some prior work, for example Ouyang et al. [47] and Taneja et al. [57], targets system-level performance programming, specifically CUDA and SIMD intrinsics. However, we are not aware of any works that address LLM code optimization for specialized hardware (i.e., not CPUs or GPUs). Hong et al. [22] show that zero-shot code generation for such languages is highly unreliable. Nonetheless, Autocomp successfully optimizes accelerator code via a combination of novel techniques.
3 The Autocomp Approach
3.1 Rationale
One naive way to generate optimized tensor accelerator code is to directly ask an LLM to rewrite the unoptimized code into its optimized counterpart. However, this approach fails for two reasons:
1.
Tensor accelerator DSLs are low-resource languages (i.e., insufficiently represented in the LLM’s training corpus), so the model produces neither semantically nor syntactically valid programs.
2.
Without guidance, the model has little notion of what optimizations to apply, or in which order.
Prior work shows that decomposing tasks, including code generation tasks, into multiple steps can improve an LLM’s ability to solve them [63, 15, 60, 22, 35]. Therefore, as shown in Fig. 6, we split our workflow into two phases: optimization plan generation and code implementation. We connect the two phases with an iterative beam search. Maintaining the best B code candidates at each step helps us explore multiple optimization trajectories in parallel. We describe our two phases and search strategy in the next section. Detailed prompts can be found in Appendix B.
In following sections, a plan refers to a natural language description of a single step of optimization and how code is transformed in that step, whereas a schedule refers to a sequence of plans that brings code from unoptimized to fully optimized.
3.2 Two-Phase Optimization
Phase 1: Optimization selection and plan generation. We prompt an LLM to select one optimization from a predefined menu of optimization options and to describe the concrete transformations required to apply it. Fig. 6 illustrates the prompt structure, which consists of the following parts:
1.
Accelerator ISA. A list of instructions in the accelerator’s ISA. We describe the semantics of each instruction in natural language, provide a specification for the accelerator’s internal memory addresses, and briefly describe the accelerator’s structure.
2.
Current Code. In the initial iteration , the original unoptimized code. When , one of the B candidates in our beam that has been selected for functional correctness and performance.
3.
Feedback. The latency of Current Code in cycles, as well as its scratchpad and accumulator utilization in kilobytes. This helps the model choose the next optimization to apply. For example, low scratchpad utilization can lead to the model suggesting a larger tile transformation. Utilization is a common feedback metric across hardware platforms that reflects how effectively we are using the accelerator’s hardware resources.
4.
Optimization Menu. A list of high-level optimizations, for instance, loop unrolling, reordering, fusion, tiling, and double buffering. Note that only the names of the optimization are included; we rely on model to generate the implementation details for the selected optimization. The full list of optimizations is in Appendix B.
5.
Instruction. A simple natural-language directive to “select exactly one optimization from the menu and output a corresponding transformation plan.”
6.
Rules. A set of constraints regarding the eventual code to be generated (see Appendix B).
7.
Beam Search Iteration. We provide the current iteration number of the beam search, as well as the total number of iterations. This helps guide optimization selection, as some optimizations are more relevant towards the beginning of search (e.g., loop splitting), whereas some are more relevant towards the end of search (e.g., loop unrolling).
At each planning iteration, we sample independent plans, seeding the search with multiple diverse optimization trajectories that can be evaluated in parallel.
Phase 2: Optimized code generation. Once we generate the candidate plans, for each plan we prompt the LLM to apply the transformations in the plan to generate a new, semantically equivalent code. In Fig. 6, we show the structure of our code generation prompt, which contains four parts:
1.
Accelerator ISA. Same as in plan generation.
2.
Current Code. Same as in plan generation.
3.
Generated Plan. The specific optimization plan generated for this code in Phase 1.
4.
In-Context Learning (ICL) Example. In cases where the optimization plan contains the string ‘‘tiling’’, we provide an example of code (from a different workload, and with variable names anonymized) before and after changing one tiling factor. Inserted as tiling is a key optimization that requires modifications across the program, making it challenging to implement.
5.
Instruction. A simple natural-language instruction to “apply the above plan and output optimized accelerator code that is functionally equivalent to the current code.”
6.
Rules. Same as in plan generation.
We sample independent code candidates for each plan because generating low-resource accelerator code is challenging, and also because our task also requires applying nontrivial transformations to the code. Sampling multiple candidates helps improve the robustness of our search process.
3.3 Beam Search
We integrate our two-phase optimization inside an iterative beam search of width . Beam search allows us to efficiently explore several optimization trajectories in parallel. Since our code mostly consists of loop nests instead of sequential code, we find that merging candidates as in prior work [36, 4, 38, 37, 16, 51] is not suitable to tensor accelerator code. As illustrated in Fig. 6, candidates from the code generation step enter the beam only if they satisfy the criteria correctness and performance:
1.
Correctness. After each code generation step, every candidate is compiled and run against our functional test suite. Each input variable is initialized with random values and after running, the candidate’s output is compared to that of a reference implementation. We first filter candidates by compiling and testing 5 times via functional hardware simulation, and up to 20 times in cycle-accurate simulation, which functionally matches real hardware at a cycle-level granularity.
2.
Performance. We measure the latency of functionally correct candidates via cycle-accurate simulation. A candidate is retained only if it improves upon the parent from which it was derived.
Of the functionally correct candidates, we keep the best (lowest latency) to seed the next iteration of beam search. Beam width is a hyperparameter, and empirically we found width the sweet spot balancing search quality and time trade-off. We run this loop for a fixed budget of iterations.
3.4 Increasing Plan and Code Diversity
We use the following two techniques to boost the diversity in plan (and in the case of LLM ensembling, code) generation and prevent the model from repeatedly selecting the same optimization:
Optimization Menu Dropout. Inspired by methods for preventing neural networks from overfitting [56], we implement dropout in our optimization menu. Each time a plan candidate is generated, each menu option in the list of optimizations has a chance to be dropped.
LLM Ensembling. Ensembling LLMs is known to improve diversity and quality of results [28]. To further increase the diversity of generated plans and code, in each case where multiple candidates are sampled, we divide these requests between different LLMs.
We ablate these techniques, along with other components of Autocomp, in Appendix A.
3.5 Schedule Reuse
Running Autocomp’s search starting with unoptimized code results in optimized code that outperforms all prior approaches, as we will discuss in Sec. 4. However, using this approach with every new software workload can be costly, as Autocomp involves multiple LLM invocations and hardware simulations. A natural question, then, is whether the schedules discovered for one workload can be used to accelerate the optimization of others. We draw inspiration from traditional optimized libraries like BLAS [3], where hand-tuned schedules are reused across GEMM shapes, and extend Autocomp with schedule reuse.
To do so, we first record the best known schedule for a particular tensor operation. Then, during planning for new GEMMs with the same aspect ratios or with two shared dimensions, rather than exploring the full menu, we prompt the LLMs to select specifically the menu options used in our recorded schedule, one at a time. As we are not exploring the full menu, we can use a smaller beam width and sample count, reducing both LLM calls and search time. After completing this lightweight search, we take the best-performing code so far and further refine it by invoking the full Autocomp search for a small number of iterations. This resembles the classic exploration-exploitation trade-off in optimization: by reusing a schedule we exploit a known high-quality schedule and avoid the initial exploration cost for a new workload.
4 Evaluating Autocomp Without Schedule Reuse
We evaluate the effectiveness of Autocomp on three distinct types of workloads 111Benchmarks uploaded to https://drive.google.com/drive/folders/1zK34x0tnzbQhcPjABxW-5uFbAKfoLGp6: 1) matrix multiplication (GEMM) derived from ResNet-50, 2) convolution derived from ResNet-50, and 3) robotics code used for model-predictive control. For all experiments, we ensemble OpenAI’s o3-mini and gpt-4o (via the OpenAI API Platform) for both phases, with temperature 1.0. Menu options are dropped out with 70% probability.
4.1 Hardware Platform
We use Gemmini [17] to generate two different accelerators for evaluation. Gemmini is an accelerator generator that can generate systolic array- and vector-style tensor accelerators with a wide range of data types and sizes. Gemmini is ideal for evaluating Autocomp as it: 1) generates accelerators that deliver performance comparable to commercial ones, 2) is open-source, enabling instantiation of different accelerator instances, user modifications to the software toolchain, and extraction of fine-grained performance feedback, and 3) supports fast and cycle-accurate hardware simulation via FireSim [31]. Like other accelerators, its low-resource nature eliminates data contamination and makes directly prompting LLMs challenging. We used AWS EC2 F1 instances and local AMD Alveo U250 FPGAs to run FireSim.
For the GEMM and convolution benchmarks in Secs. 4.3 and 4.4, we use Gemmini to generate a 1616 systolic array on 8-bit integer data type (accumulating in 32-bit), 256 KB scratchpad, and 64 KB accumulator, the same platform used by Ikarashi et al. [26]. For the fine-grained linear algebra benchmarks in Sec. 4.5, we generate a 32-bit floating point accelerator with a 44 systolic array, and the same scratchpad and accumulator sizes, as used by Dong et al. [14]. This information is provided in the Instruction portion of our planning prompt.
Approach | Performance | GEMM/conv size-agnostic | Supports workloads other than GEMM/conv | No power/ area cost |
Generic library (e.g. Gemmini SW Lib, Exo Unopt) | Low | ✓ | ✗ | ✓ |
Hand tuning (e.g. Exo Opt) | Medium/High | ✗ | ✓ | ✓ |
Hardware FSM | High | ✓ | ✗ | ✗ |
[…]
4.2 Baselines
For the first two workload types, we compare Autocomp with four baselines, characterized in Table 1:
1.
Gemmini’s Software Library. Gemmini ships with a software library that uses heuristics to tile and run GEMMs and convolutions on generated accelerators. As loop ordering is fixed and loop bounds, addresses, and indices must be computed at runtime, this implementation incurs significant software overhead and cannot fully utilize hardware resources.
2.
Exo Unoptimized. Exo [26] is a DSL for tensor computation and scheduling. It comes with a basic compiler that emits functional GEMM or convolution code that can be executed on accelerators generated by Gemmini. However, without benchmark-specific optimization, performance is highly suboptimal, as hardware resources such as local memories tend to be underutilized.
3.
Exo Optimized. In Ikarashi et al. [26], Exo’s and Gemmini’s developers spent significant effort manually writing and hand-tuning benchmark- and accelerator-specific schedules for each of the GEMM and convolution sizes in Figs. 8 and 8. This is the previous best known software implementation for these benchmarks.
4.
Hardware FSM. Gemmini can generate accelerators with specialized hardware units for two coarse-grained operations: GEMM and convolution. These hardware units, implemented as finite state machines (FSMs), encode control sequences for each of these operations in hardware. If tuned correctly, the FSMs can exceed the theoretical maximum performance of any software-based implementation, as hardware is inherently parallel. However, this is accomplished at the cost of scheduling flexibility, as well as increased area, power, and hardware complexity. We use the hardware FSM as a reference for the highest achievable performance, but do not expect to exceed its performance for these GEMM/convolution benchmarks as its compute utilization exceeds 90% for all but one benchmark (82% on 12544x256x64). Despite this, Autocomp-generated code approach hardware FSM performance for GEMM/convolution, and as seen in Sec. 4.5, even exceeds hardware FSM performance in end-to-end application performance thanks to Autocomp’s greater scheduling flexibility.
For GEMM and convolution, we use Exo Unoptimized, which contains statically pre-computed loops, addresses, and indices, as Autocomp’s starting point. This simplifies code generation and allows us to directly compare the effectiveness of Autocomp to hand-optimization (i.e., Exo Optimized code).
The third workload type, robotics control code, is a multi-kernel application containing sequences of element-wise operations and matrix-vector multiplications. As this is not directly supported by Exo, we compare to an unoptimized software implementation ported to accelerator code by Dong et al. [14], and an extensively hand-optimized hardware FSM-based implementation written by an expert.
4.3 Matrix multiplication (GEMM)
We run Autocomp on a set of GEMM benchmarks selected by Ikarashi et al. [26] from ResNet-50 [19] for diversity in size and shape. We run search with beam size , plans per element in the beam, code candidates per plan, and iterations. This takes around 5 hours to run.
Fig. 8 shows that Autocomp significantly outperforms even extensively hand-optimized code (Exo Opt) by a geomean of 1.4, Exo Unoptimized code (the starting point of Autocomp’s search) by 2.9, and Gemmini’s software library by 5.6. Autocomp is consistent: its generated code always achieves at least 85% of the hardware FSM’s utilization (and 91% on average).
Autocomp especially outperforms prior implementations thanks to extensive exploration of software pipelining and double-buffering, which allows better overlapping of data movement and computation, for example by double-buffering both the scratchpad and accumulator. In many cases, Autocomp’s exploration also leads to different tiling and loop ordering choices than hand-optimized code, reducing data movement. We qualitatively analyze Autocomp-generated GEMM code in detail in Appendix C.
4.4 Convolution
We also optimize convolution benchmarks from ResNet-50 via the same process. Compared to the GEMM benchmarks, this code contains more loops and is more complex. In this case, we run beam search with beam size , plans, and code candidates, for iterations, which takes about 7 hours.
Compared to GEMM, convolution provides less room for improvement over both the Gemmini software library and Ikarashi et al. [26]’s implementation, as even the hardware FSM only achieves a 3.3 geomean speedup over the software library, compared to 6.1 for GEMM. This is because on average the Gemmini software library achieves 28% of the theoretical maximum compute utilization, compared to 16% for GEMM. As discussed by Genc et al. [17], at the tensor sizes in ResNet-50, convolutions have greater arithmetic intensity than GEMMs, making them less memory-bound and causing the Gemmini software library’s suboptimal data orchestration to be less impactful.
Nonetheless, as shown in Fig. 8, Autocomp still exceeds the previous best known hand-optimized software ISA-based implementation (Exo Opt) by up to 1.2 and by a geomean of 1.1, via similar strategies as for GEMM. It also outperforms Exo Unoptimized code by 2.9, and Gemmini’s software library by 2.6, and in all cases achieves at least 78% of the hardware FSM’s utilization.
4.5 Fine-Grained Linear Algebra
Finally, we optimize fine-grained linear algebra benchmarks from the TinyMPC model-predictive control library [40], specifically the forward and backward passes of the primal update step. These benchmarks contain sequences of floating-point matrix-vector multiplications, interleaved with element-wise addition and subtraction. The inclusion of CPU-accelerator dependencies, low reuse, and a high ratio of data movement to computation leads to low accelerator utilization and makes this code challenging to optimize.
We compare Autocomp-generated code against Dong et al. [14]’s unoptimized software-based implementation on a 44 FP32 accelerator. For this work, we additionally had an expert hand-tune a hardware FSM-based implementation. The unoptimized software-based implementation is used as the starting point for search, and we use the same search parameters as for convolution, except with iterations, which takes about 12 hours. However, some of the optimization menu options are different from those used for GEMM/convolution (see Appendix B). As shown in Fig. 9, Autocomp outperforms even the expert-optimized hardware FSM implementation on the forward pass (by 1.6), and across benchmarks speeds up unoptimized code by a geomean of 2.7.
To outperform the hardware FSM implementation, Autocomp harnesses the flexibility of software-based implementation. It optimizes the code by hoisting data loads shared between kernels (reducing data movement beyond what is possible for the hardware FSM implementation), as well as utilizing fine-grained software pipelining and eliminating blocking operations where possible. This experiment highlights Autocomp’s adaptability: we optimize a new benchmark, running on an accelerator with a new size and data type, with highly different performance characteristics from previous experiments, by updating the Accelerator ISA and changing a few lines in the Optimization Menu and Instruction sections of the prompt.
Category | Base Benchmark | Reuse Targets |
Square | 1024×1024×1024 | 512×512×512, 256×256×256 |
Column-dominant | 12544×256x64 | 6272×256x64, 12544×128x64 |
Row-dominant | 128×1024×1024 | 64×1024×1024, 32×1024×1024 |
5 Improving Sample Efficiency With Schedule Reuse
In this experiment, we illustrate that optimization schedules generated by Autocomp for one tensor operation can be effectively reused to accelerate the optimization of similar tensor operations (as described in Sec. 3.5). For this experiment, we generate additional benchmarks structurally similar (either same aspect ratio or two shared dimensions) to those in Sec. 4.3. For each new benchmark, in the plan reuse phase, we search with beam size , plan candidates, and code candidates. Once the reuse sequence is exhausted, we switch to a full Autocomp search with beam size , plan candidates, and code candidates.
We consider GEMMs from the following categories: 1) square, 2) column-dominant and, 3) row-dominant, with dimensions specified in Table 2. The “base” benchmark is optimized using full search, i.e., without any schedule reuse, and its Autocomp generated schedule is then reused to optimize the other two benchmarks in its category.
We include the full optimization traces for Autocomp with and without reuse for two example benchmarks: 512×512×512 (square) and 12544x128x64 (column-dominant). As shown in Fig. 10, with reuse, Autocomp consistently achieves better speedups with the same number of samples.
For the remaining benchmarks, we perform an iso-samples comparison: we fix the total number of LLM calls and compare reuse to full search. In Fig. 10, the red vertical lines compare the iso-sample performance for our two examples. Across all reuse targets in Table 2, with a budget of 100 samples (10% of the calls used by full search), Autocomp achieves geomean speedups over Gemmini’s software library of with reuse, compared to without reuse. With a 200 call budget, we observe speedup with reuse, compared to without reuse. This demonstrates Autocomp schedules are generalizable and can help reduce the search cost when optimizing similar benchmarks.
6 Conclusion
In this paper, we demonstrate how to construct an LLM-based flow to automatically optimize low-resource accelerator code at superhuman levels. Autocomp outperforms the previous best known software implementation by generating code that is significantly faster on real-world benchmarks. By reducing the amount of manual effort needed to build compilers, we show the effectiveness of Autocomp’s approach and its potential to serve as a key component in the accelerator design process.
While Autocomp already delivers impressive performance across all benchmarks, in future work there is scope for broader plan reuse for greater sample efficiency. Currently, we only reuse individual schedules; we can go further by extracting common patterns across schedules and summarizing them as “meta-schedules” which explain how to schedule and can be applied more broadly.
Appendix A Ablation Studies
Experiment | 12544x64x256 GEMM Speedup | 4x3x14x256x256 Conv Speedup |
Baseline (Exo Unopt) | 1.67 | 0.91 |
No Accelerator ISA | 3.11 | 2.51 |
No Optimization Menu | 2.34 | 0.97 |
No Optimization Menu Dropout | 4.72 | 2.30 |
No LLM Ensemble (o3-mini only) | 4.67 | 2.08 |
No Hardware Perf Feedback | 4.91 | 2.61 |
LLM Selection (DeepSeek-R1) | 4.89 | 2.25 |
Autocomp | 5.53 | 2.72 |
In this section, we ablate various features of Autocomp to investigate their effect on optimization performance. We focus on two specific benchmarks from Sec. 4—our 12544x64x256 GEMM and our 4x3x14x256x256 convolution—to isolate the effects of these ablations while limiting the cost of running this extensive exploration.
A.1 Accelerator ISA
We find that for GEMM and convolution code, removing the ISA significantly deteriorates performance. Still, Autocomp is able to improve over the original code by a notable margin even without the ISA, given that all its other features are still enabled (see Table 3). This is because we inherently provide an example of accelerator ISA code at each step via the current code, so the model is able to infer some properties of the accelerator ISA. In addition, many of the nested loop optimizations for the GEMM and convolution workloads are well-understood transformations that operate on the C-syntax loops, addresses, and indices in the code, which resemble general-purpose programming, rather than using accelerator-specific constructs such as configuration instructions. However, full Autocomp performance is not matched as the proportion of functionally correct responses is lower, and instruction-level optimizations cannot easily be identified. For example, the first-compute handling optimization in Appendix C.1’s GEMM example and the negative-scaled bias loading in Appendix C.2’s fine-grained linear algebra example would not have been identified without knowledge of the ISA. Overall, we find that the accelerator ISA is an important part of Autocomp’s prompts.
A.2 Optimization Menu
We ablate the optimization menu by removing the menu in the planning prompt, and instead simply asking the model to select one optimization and generate a plan. From this experiment, we find that adding domain knowledge and optimization diversity via the optimization menu is essential to Autocomp. As shown in Table 3, optimization performance completely deteriorates without the optimization menu. Qualitatively, without the optimization menu, we find that the models tend to repeat similar optimizations, with significantly less diversity and relevance in the generated optimization plans.
A.3 Optimization Menu Dropout
“Dropout” for optimization menu options is a key contribution of this work that increases the diversity of generated optimization plans. Table 3 shows that menu dropout has a significant effect on performance. Qualitatively, we find that without dropout, models tend to be biased towards a limited set of menu options, a limitation which can be resolved via menu dropout.
A.4 Hardware Performance Feedback
As discussed in Sec. 3.2, during plan generation, we include the latency, scratchpad utilization, and accumulator utilization of the original code. Table 3 shows that this component is helpful, but in some cases its effects may be limited. This is because the options listed in the optimization menu already capture some of the metrics measured in our performance feedback, for example the menu options which suggest using larger tile sizes. Hardware feedback such as scratchpad and accumulator utilization only serves to augment elaboration of these menu options by providing exact measurements.
A.5 LLM Ensembling
Splitting requests between LLMs in an ensemble also encourages diversity of generated plans and code. Qualitatively, we find that the responses, especially during the planning phase, generated by different LLMs differ substantially. Our experiments in Table 3 show that using individual models, such o3-mini or DeepSeek-R1, on their own results in significantly lower performance.
A.6 LLM Selection
In Sec. 4, we use an ensemble of gpt-4o and o3-mini for our search. To demonstrate that Autocomp does not depend on a particular family of models, we run Autocomp with DeepSeek-R1 [12] on the same benchmarks used for other ablation experiments above. We use DeepSeek-R1 for both the planning and code generation phases and keep the search parameters identical to those used for matrix multiplication and convolution in Sec. 4. As shown in Table 3, Autocomp with DeepSeek-R1 is able to optimize both GEMM and convolution, achieving substantial speed-ups over the unoptimized code. These experiments do not use LLM ensembling. Similarly to the LLM ensembling ablation study above, gains are slightly smaller than when all techniques are applied. Nonetheless, this demonstrates that Autocomp is efficient and flexible across different LLMs.
Appendix B Prompts
Appendix C Code Examples
In this section, we discuss in greater depth what optimizations Autocomp applies in our evaluations and how Autocomp is able to achieve significantly better performance than hand-optimized code.
C.1 12544x64x256 GEMM
Fig. 17 contains the unoptimized Exo-generated code, used as the starting point for search. Fig. 18 contains the code generated by Exo after hand-optimization by Ikarashi et al. [26]. Figs. 19, 20 and 21 contain the result of Autocomp optimization on Exo Unoptimized code. While the code is substantially transformed from the original code, some aspects remain the same. For example, in this case the configuration instructions and loop ordering remain largely the same.
Of course, many optimizations have been applied to the code. We briefly summarize the optimization menu options selected and plans generated during the optimization process for this code. We also include the speedup after each respective optimization.
1.
1.67: initial speedup of Exo Unoptimized code over Gemmini’s software library before any optimization.
2.
1.93 after “hoist redundant operations out of loops”. This plan hoists constants like tile_dim = 16 and loop-invariant expressions like ko * 64 and k * 16 out of inner loops. These precomputed values are reused inside Gemmini ops in each iterations, so we should reduce the number of times they must be calculated.
3.
1.95 after “double buffering”. This plan defines two buffer regions for matrices A and B in the scratchpad. A buffer_toggle flag is introduced to alternate between these buffers each iteration. All mvin, preload, and compute instructions are updated to use the active buffer based on the toggle. Data loading for the next iteration is scheduled earlier to overlap with current computation. Address calculations are adjusted to include buffer offsets accordingly.
4.
2.15 after “pipeline operations to better overlap computation and data movement”. Moves mvin2 (A tile load) to immediately after compute_preloaded in the ko loop to overlap A prefetch with current compute. Moves mvin3 (B tile load) earlier in the k loop, before the next compute, to overlap B prefetch with current compute.
5.
3.13 after “load data to the scratchpad across outer loop iterations and use if statements to prevent redundant loads on loops inner to those”. Adds a one-time load of the entire B matrix into a new scratchpad region (specifically, new_B_base = 8192) before the main i loop. Replaces repeated mvin3 calls in inner loops with offset calculations into this preloaded B region.
6.
3.54 after “move more data to the scratchpad in a more outer loop to increase data reuse”. Before the reduction (k) loop, load the entire 16x256 A-tile into scratchpad using four mvin2 calls to a new A_tile_base region. Remove the double-buffering (a_toggle) for A since the full tile is now resident. Replace nested ko and k loops with a single loop over 16 segments, computing each 16x16 A sub-tile address from A_tile_base. Adjust B tile preload and compute_preloaded calls accordingly, using a simpler tile index.
7.
4.87 after “double buffering”. Restores double buffering by reserving two scratchpad regions: A_tile_base0 and A_tile_base1, each holding one full 16x256 A-tile. In the outer i loop, alternate buffers using i % 2 to select current_buffer and next_buffer. In each iteration (except last), while computing with current_buffer, issue mvin2 to load the next A-tile into next_buffer. In the inner compute loops, use current_buffer for A-tile addresses. After compute and mvout, the next iteration uses the preloaded data.
8.
5.21 after “double buffering”. Double-buffers accumulator by allocating two accumulator regions: acc_base0 and acc_base1. In each i iteration, compute into cur_acc_base and mvout from prev_acc_base (except on first iteration), and swap cur_acc_base and prev_acc_base at the end of the loop.
9.
5.23 after “loop unrolling”. Unrolls the innermost loop (j_in_o) by a factor of 4.
10.
5.53 after “fuse loops”. “Fuses” loops by eliminating the loop over j_in_o where we mvin 0s to the accumulator. Instead, use the ability of preload to overwrite the values in the accumulator rather than accumulating, when beginning a new partial sum.
From this example, we observe that a diverse set of optimizations is selected, and that speedups are distributed throughout the optimization process rather than concentrated in just one or two steps, showing the importance of a well-designed iterative search process. From here, we summarize the differences between Autocomp-generated code and the previous best code (Exo Opt):
Tiling. The Exo Opt code loads 128256 tiles of A, whereas the Autocomp-generated code loads 32256 tiles (divided into two 16256 tiles) of A. While this means there is less reuse for the Autocomp-generated code, there is also less overhead needed for compute instructions to wait for each A tile to be loaded to the scratchpad. In combination with the rest of the optimizations applied by Autocomp, this leads to improved performance.
Double-buffering. In the Autocomp-generated code, we see that both the scratchpad and accumulator are explicitly double-buffered. In the schedule, we can see that double buffering is applied 3 times. Initially (in step 3), both the A and B matrices (where matrix multiplication is represented as ABC), are double buffered in the scratchpad. However, after steps 5 and 6, B and A (respectively) are no longer double-buffered as larger tiles are loaded before beginning computation. The accumulator is double-buffered in step 8, resulting in the code below. The Exo Opt code relies on the accelerator’s out-of-order execution to handle executing mvin and mvout instructions without dependencies, ahead of the order in which they are issued.
Software pipelining. The Autocomp-generated code explicitly issues A mvin instructions before they are needed for computation, whereas as above the Exo Opt code relies on hardware to handle overlapping of data movement and compute. Also, the Autocomp-generated code explictly issues all B mvin instructions at the beginning of the program, whereas the Exo Opt code interleaves these instructions with computation (but still loads the entire B matrix to the scratchpad, once overall). This does not have a significant impact on performance, but LLM-generated code is naturally biased towards such an implementation due to its simplicity.
First-compute handling. The Autocomp-generated code utilizes the ability of compute instructions to overwrite the accumulator, whereas Exo Opt code explicitly issues mvin instructions to zero out the accumulator before beginning computation on a tile.
Arithmetic simplification. Arithmetic on constants is fully simplified and handled inside shared variables wherever possible in the Autocomp-generated code, reducing the overhead of non-accelerator instructions.
Overall, we find that compared to the Exo Opt code, Autocomp-generated code applies more techniques to minimize the amount of CPU overhead during execution. The smaller tiles it uses, in combination with its more explicit application of double-buffering and software pipelining, results in highly tuned, fine-grained overlapping of data movement and computation and a very high level of performance.
C.1.1 Unoptimized Code Example
C.1.2 Exo Optimized Code Example
C.1.3 Autocomp-Optimized Code Example
C.2 TinyMPC Primal Update Forward Pass (Fine-Grained Linear Algebra)
The forward pass of the TinyMPC primal update step computes the following operations:
Where is a 12x12 matrix, is a 12x4 matrix, is a 4x12 matrix, is an NHORIZONx12 matrix (where individual columns are accessed here via indexing), is an NHORIZONx4 matrix, and is an NHORIZONx4 matrix. , , , and the 0th column of are inputs, and is the output. The 1st to (NHORIZON-1)th column of are intermediate values computed over the course of the benchmark.
This process is repeated until a time horizon, defined as NHORIZON in our code and set to 5 for our evaluations. Note that is allocated as an (NHORIZON+1)x12 matrix in our code since the unoptimized code accesses up to the (NHORIZON)th column.
Autocomp generates the code in Fig. 25, optimized over several steps from the starting code in Fig. 17. The following optimizations are applied, with the following speedups after each optimization:
1.
1: in this case we treat the unoptimized software as the baseline for speedup, so by definition its speedup is 1.
2.
1.07 after “hoist redundant operations out of loops”. Hoists the mvin calls for the constant matrices Kinf, Adyn, and Bdyn above the NHORIZON loop and executes them once rather than in every iteration. Any associated config_ex and config_ld calls are also moved outside the loop if needed. The compute calls use the same scratchpad addresses and may set B_spad_addr = 0xffffffff to indicate that the weights are already loaded.
3.
1.13 after “loop reordering”. This plan claims to move the configuration and mvin instructions for Kinf, Adyn, and Bdyn before the NHORIZON loop, but this has already been handled in the previous step. In reality, only some configuration instructions that were unnecessarily left behind by the previous step are hoisted.
4.
2.00 after “move CPU-based computation to the accelerator”. Replaces CPU-based element-wise negation and addition with equivalent Gemmini compute instructions. Specifically, when x_i and d_i are loaded, a scaling factor of -1 is used to negate it in the scratchpad. The product of Kinf and x_i is kept in the accumulator and the negated d_i is multiplied by a dummy 0 vector in order accumulate it on top of this product, enabling data to be kept resident in the accumulator rather than moving it back and forth between the CPU and accelerator memory.
5.
2.12 after “hoisting redundant operations out of loops”. Identifies a few config_ld instructions that are still redundant and removes them from the loop.
6.
2.66 after “loop unrolling”. Changes the outer loop to increment by 2 instead of 1 and duplicate the loop body so each iteration computes two time steps: first u[i] and x[i+1], then u[i+1] and x[i+2]. Also merges fence calls at the end of the unrolled loop body if possible. The implementation very aggressively removes fence instructions, which is actually correct as reuse of the same accelerator memory instructions by subsequent mvin and mvout instructions means that dependencies can be handled internally to the accelerator, rather than via synchronization of both the CPU and accelerator.
7.
2.95 after “loop fusion”. Undoes the loop unrolling from the previous step, but keeps the reduced fence instructions. Additionally, this eliminates the unnecessary calculation of x[NHORIZON] during execution, which saves cycles. This optimization makes sense since it is likely the reduced number of fence instructions that improved performance in step 6, rather than the loop unrolling (which usually provides limited benefit, if any).
We further discuss the similarities and differences between the hand-optimized hardware FSM-based implementation in Fig. 24 and the Autocomp-optimized code:
Data orchestration. Due to the coarse-grained nature of the hardware FSM operations and the fact that data movement between the accelerator and main memory is handled within hardware, we are not able to hoist shared data loads of the Kinf, Adyn, and Bdyn matrices out of the loop. This is the main advantage the Autocomp-generated software ISA-based implementation has over the hardware FSM-based implementation.
Operator fusion. Both implementations are able to handle all computation on the accelerator, but notably the hardware FSM-based implementation fuses the addition of d[i] into the bias when multiplying Kinf and x[i] (then the whole result is negated while storing to main memory), whereas the addition of d[i] is handled as a separate compute instruction in the Autocomp-generated code. So, the hardware FSM-based implementation actually has an advantage in this regard and the Autocomp-generated code has further room for improvement. Both implementations make use of negative scaling of loaded data (via config_ld instructions) in order to handle subtraction.
Fence instructions. The Autocomp-generated code is able to remove CPU-based fence instructions and instead handle dependencies within the accelerator, whereas the hardware FSM-based code is forced to place a fence after each matrix multiplication, as accumulated results must be moved back to main memory and loaded to the scratchpad for the next operation.
Configuration overhead. While they do not have as much overhead as fence instructions, configuration instructions can also be blocking. The hand-optimized code hoists configuration instructions out of the loop where possible, but since different matrix sizes must be loaded inside the loop, configuration instructions cannot be completely eliminated, giving Autocomp’s code the advantage in this aspect.
Dead code elimination. Both implementations eliminate the extra computation of x[NHORIZON] that is present in the unoptimized code.
Overall, we find that Autocomp identifies all major optimization opportunities available in the code, with the exception of handling subtraction of d[i] slightly suboptimally. Qualitatively, optimizing this code by hand can be difficult due to a lack of readability and the difficulty of debugging low-level accelerator code. The Autocomp-generated code is well-commented and the sequence of optimizations applied can be easily understood. This will be helpful for further optimization of this benchmark or future optimization of other benchmarks, via methods like the schedule reuse demonstrated in Sec. 5.