### v5

Abstract
Hardware accelerators, especially those designed for tensor processing, have become ubiquitous in today’s computing landscape. However, even with significant efforts in building compilers, programming these tensor accelerators remains challenging, leaving much of their potential underutilized. Recently, large language models (LLMs), trained on large amounts of code, have shown significant promise in code generation and optimization tasks, but generating low-resource languages, such as specialized tensor accelerator code still poses a significant challenge. We tackle this challenge with Autocomp, an approach that empowers accelerator programmers to leverage domain knowledge and hardware feedback to optimize code via an automated LLM-driven search. We accomplish this by: 1) formulating each optimization pass as a structured two-phase prompt, divided into planning and code generation phases, 2) inserting domain knowledge during planning via a concise and adaptable optimization menu, and 3) integrating correctness and performance metrics from hardware as feedback at each search iteration. Across three distinct hardware platforms, we demonstrate that Autocomp-optimized code runs 5.6 faster than the vendor-provided library (Gemmini), outperforms expert-level hand-tuned code by 1.9 (AWS Trainium), and achieves 3.8 higher performance than a machine learning-based cost model for GPUs (NVIDIA L40S). Additionally, we demonstrate that optimization schedules generated from Autocomp can be reused across similar tensor operations, improving speedups by up to 24% under a fixed sample budget.
Preliminary work. Under review by the Machine Learning and Systems (MLSys) Conference. Do not distribute.
1 Introduction
Hardware accelerators nickolls2008cuda; tpu-isca2016 have become a critical driving force for the recent breakthroughs alexnet; resnet; transformer; chatgpt in machine learning. They provide orders-of-magnitude improvements in performance and energy efficiency in running deep neural networks (DNNs), and this has led to an increasing number of accelerators for tensor processing in recent years ane; nvdla; lauterbach2021cerebras; intelamx; armsme. However, extracting that performance requires writing high-performance accelerator code, which is time-consuming and requires a deep understanding of the underlying hardware.
To address this challenge, various compilers and domain-specific languages (DSLs) have appeared. For deep learning applications, compilers such as XLA, TVM, and Triton generate high-performance code, but they only support a few hardware backends, particularly CPUs and GPUs xla; chen2018tvm. Unfortunately, adapting compilers and DSLs to new hardware platforms with vendor-specific instruction set architectures (ISAs) and implementation-specific dataflow patterns requires significant engineering effort. In fact, software alone comprises 40-50% of the development cost for new hardware reuters2025openaichipcost, even before considering the effort needed for end users to write and debug software for a newly developed chip. Prior work in DSLs like Halide and Exo jrk2013halide; exo targets accelerators by providing primitives that make it easier to express tensor computation, but the onus of optimizing code written in such DSLs still lies on the accelerator programmer.
Even once a compiler exists, generating performant code runs into the classical “scheduling” problem, i.e., deciding which optimizations to apply and in what order. For general-purpose backends (CPUs and GPUs), these passes have been iteratively developed and refined over many years through a combination of experts and auto-tuning frameworks. Recent work has gone further in exploring data-driven approaches such as supervised learning zheng2021tenset, reinforcement learning cummins2022compilergym, and even LLM training cummins2023compiler; metallmcompiler to tackle the combinatorial explosion of pass sequences. While these data-driven approaches have shown promise, they depend on vast amounts of performance data to train, which is painfully scarce for domain-specific hardware accelerators.
In this paper, we present Autocomp, which solves the problems with prior approaches with an iterative LLM-driven search framework to optimize accelerator code. Unlike previous compilers, Autocomp can adapt to new hardware platforms and ISAs by simply changing prompts. Unlike existing tensor DSLs, Autocomp automatically generates optimized code without manual tuning. And unlike data-driven approaches targeting CPUs and GPUs, Autocomp requires no model training, instead leveraging LLM in-context reasoning and pretrained knowledge of common optimizations.
In each iteration, Autocomp first plans by choosing an optimization from a predefined menu, i.e., a list of common hardware accelerator optimizations like tiling and unrolling, then applies the optimization to generate optimized DSL code. The generated candidates are validated for correctness and benchmarked on the accelerator to collect performance metrics, providing feedback for the next iteration of search. By encoding DSL syntax, optimization rules, and performance feedback concisely in a prompt, Autocomp guides the LLM to generate optimized accelerator code.
A key advantage of this prompt-driven approach is its inherent portability. Whereas traditional compilers require significant engineering effort to target a new hardware platform, Autocomp can be retargeted by modifying its structured prompt. This dramatically lowers the barrier to entry for optimizing code for new low-resource accelerators. We demonstrate this in our experiments by applying Autocomp to three distinct hardware backends.
In our evaluation, we find that Autocomp significantly outperforms prior approaches across hardware platforms. On Gemmini, Autocomp speeds up GEMMs by 5.6 compared to the vendor-provided library and 1.4 compared to expert-level hand-tuned code, surpassing the prior best known implementations with less human effort. On AWS Trainium, Autocomp reduces the runtime of a wide range of hand-optimized tensor operators by an average of 1.9. On an NVIDIA L40S GPU, Autocomp generates KernelBench kernels that run 2 faster than PyTorch and 3.8 faster than TVM MetaSchedule. Moreover, we show that Autocomp’s schedules can be reused as guidance when scheduling similar tensor operations, alleviating the cost of scheduling new code and delivering up to 24% greater speedups under a fixed sample budget.
In summary, we make the following contributions:
1.
We present Autocomp, the first LLM-driven code optimization approach for low-resource tensor accelerator code generation.
2.
Autocomp is a portable optimization framework that dramatically lowers the engineering effort to target new hardware. Our search incorporates domain knowledge, hardware feedback on correctness and performance, and novel strategies for response diversity to automatically generate performant code.
3.
Autocomp-generated code significantly outperforms expert hand-optimized code across a wide range of workloads and across different tensor accelerators.
4.
We illustrate that schedules generated by Autocomp can be reused to optimize similar tensor operations, reducing search cost and demonstrating the a posteriori usefulness of Autocomp-generated schedules beyond pure performance.
5.
Our implementation and prompts are fully open-source111https://github.com/ucb-bar/autocomp.
2 Background
2.1 Code Optimization for Tensor Accelerators
Programming tensor accelerators differs greatly from programming general-purpose CPUs. Tensor accelerators, depicted in Fig.˜1, generally focus on the efficient execution of fixed-size (e.g., 1616) matrix multiplication instructions, as shown in Fig.˜2. Rather than trying to reduce the number or type of these instructions, which is often fixed, software optimization focuses on other aspects, such as:
Minimizing data movement between main memory and smaller accelerator-local memories (in Fig.˜1, the scratchpad and accumulator).
Setting configuration state for computation and data movement.
Scheduling or reordering operations to maximally overlap computation and data movement.
Code transformations that enable these optimizations range from low-level changes like arithmetic simplification or instruction selection, to higher-level changes like loop tiling, hoisting (Fig.˜3), or software pipelining (Fig.˜4). These high-level changes, while improving performance, require loop nests, pointers, and indices to be modified in multiple locations, making them challenging to implement, especially in a low-resource DSL.
Prior work has explored some of this optimization space. For example, performance models like Timeloop parashar2019timeloop and MAESTRO kwon2020maestro use high-level hardware architectural models and software abstractions to represent tensor accelerators and their workloads. Much recent work has sought to automatically explore this space, using methods such as machine learning dosa; vaesa; sakhuja2024polaris, linear programming huang2021cosa, black-box optimization zhang2022fast; sakhuja2023spotlight, and reinforcement learning xiao2021hasco. While these abstractions capture some aspects of tensor accelerator code optimization, in particular the amount of data movement, they neglect other implementation-specific and instruction-level optimizations. In this work, LLM-based code generation allows us to directly rewrite accelerator code, expanding the search to include all potential axes of optimization.
2.2 LLM-Based Code Optimization
LLMs have been used in various code-related tasks chen2021humaneval; nijkamp2022codegen; bhatia2024llmlift; hong2025hdl2vcodetranslationdataset. For code optimization, researchers have successfully used LLMs in various paradigms, including evolutionary search lehman2022evolutionlargemodels; romeraparedes2024funsearch, retrieval-augmented generation wang2025symrtlo; lin2025ecollmdrivenefficientcode, iterative refinement using compiler feedback peng2024perfcodegenimprovingperformancellm; damani2024warpdrive, and model post-training shypula2024learning; baronio2025kevinmultiturnrlgenerating.
Some prior work, for example ouyang2025kernelbench and loopvectorizer, targets system-level performance programming, specifically CUDA and SIMD intrinsics. However, we are not aware of any works that address LLM code optimization for specialized hardware (i.e., not CPUs or GPUs), despite the growing need for effective code optimization methods. hong2024llmaidedcompilation show that zero-shot code generation for such languages is highly unreliable. Nonetheless, Autocomp successfully optimizes accelerator code via a combination of novel techniques.
3 The Autocomp Approach
3.1 Rationale
One naive way to generate optimized tensor accelerator code is to directly ask an LLM to rewrite the unoptimized code into its optimized counterpart. However, this approach fails for two reasons:
1.
Tensor accelerator DSLs are low-resource languages (i.e., insufficiently represented in the LLM’s training corpus), so the model produces neither semantically nor syntactically valid programs.
2.
Without guidance, the model has little notion of what optimizations to apply, or in which order.
Prior work shows that decomposing tasks, including code generation tasks, into multiple steps can improve an LLM’s ability to solve them gao2023pal; wang2024planningnaturallanguageimproves. Therefore, as shown in Fig.˜5, we split our workflow into two phases: optimization plan generation and code implementation. We connect the two phases with an iterative beam search. Maintaining the best B code candidates at each step helps us explore multiple optimization trajectories in parallel. We describe our two phases and search strategy in the next section. Detailed prompts are in Appendix C.
In following sections, a plan refers to a natural language description of a single step of optimization and how code is transformed in that step, whereas a schedule refers to a sequence of plans that brings code from unoptimized to fully optimized.
3.2 Plan-then-Implement: Two-Phase Optimization
Phase 1: Plan. We prompt an LLM to select one optimization from a predefined menu of optimization options and to describe the concrete transformations required to apply it. Fig.˜5 illustrates the prompt structure, which consists of the following parts:
1.
Accelerator ISA. A list of instructions in the accelerator’s ISA. We describe the semantics of each instruction in natural language, provide a specification for the accelerator’s internal memory addresses, and briefly describe the accelerator’s structure.
2.
Current Code. In the initial iteration , the original unoptimized code. When , one of the B candidates in our beam that has been selected for functional correctness and performance.
3.
Feedback. The latency of Current Code in cycles, as well as (for Gemmini) its scratchpad and accumulator utilization in kilobytes. Utilization is a common feedback metric across hardware platforms that reflects how effectively we are using the accelerator’s hardware resources, and can help the model choose the next optimization to apply. For example, low scratchpad utilization can lead to the model suggesting a larger tile transformation.
4.
Optimization Menu. A list of high-level optimizations, for instance, loop unrolling, reordering, fusion, tiling, and double buffering. Note that only names or short descriptions of the optimization are included; we rely on the model to generate the implementation details for the selected optimization. Furthermore, the menu options range in specificity and include an option to select an optimization not listed, so the model’s output is not overly constrained. The full list of optimizations for each hardware platform is in Appendix C.
5.
Instruction. A high-level description of the target accelerator, followed by a natural-language instruction to “select exactly one optimization from the menu and output a corresponding transformation plan.”
6.
Rules. A set of constraints regarding the eventual code to be generated (see Appendix C).
7.
Beam Search Iteration. The current iteration of beam search iterations. Helps guide optimization selection, as some optimizations are more relevant towards the beginning of search (e.g., loop splitting) or towards the end of search (e.g., loop unrolling).
At each planning iteration, we sample independent plans, seeding the search with multiple diverse optimization trajectories that can be evaluated in parallel.
Phase 2: Implement. Once we generate the candidate plans, for each plan we prompt the LLM to apply the transformations in the plan to generate a new, semantically equivalent code. Fig.˜5 shows the structure of our code generation prompt, which contains the following parts:
1.
Accelerator ISA. Same as in Phase 1.
2.
Current Code. Same as in Phase 1.
3.
Generated Plan. The specific optimization plan generated for this code in Phase 1.
4.
In-Context Learning (ICL) Example. Depending on the target hardware platform, we include examples of certain difficult-to-implement optimizations.
5.
Instruction. A simple natural-language instruction to “apply the above plan and output optimized accelerator code that is functionally equivalent to the current code.”
6.
Rules. Same as in Phase 1.
We sample independent code candidates for each plan to improve the robustness of our search, since generating low-resource accelerator code is challenging and our task requires applying nontrivial transformations to the code.
3.3 Beam Search
We integrate our two-phase optimization inside an iterative beam search of width . Beam search allows us to efficiently explore several optimization trajectories in parallel. Since our code mostly consists of loop nests instead of sequential code, we find that merging candidates as in prior work lehman2022evolutionlargemodels; ma2024eurekahumanlevelrewarddesign; novikov2025 is not suitable to tensor accelerator code. As illustrated in Fig.˜6, candidates from the code generation step enter the beam only if they satisfy the criteria correctness and performance:
1.
Correctness. After each code generation step, every candidate is compiled and run against our functional test suite. Each input variable is initialized with random values and after running, the candidate’s output is compared to that of a reference implementation. We first filter candidates by via a functional correctness check on the target hardware.
2.
Performance. We measure the latency of functionally correct candidates via cycle-accurate simulation or on-chip performance measurement (depending on the hardware platform). A candidate is retained only if it improves upon the parent from which it was derived.
Of the functionally correct candidates, we keep the best (lowest latency) to seed the next iteration of beam search. Empirically, we find beam width to be a sweet spot that balances search quality and time trade-off. We run this loop for a fixed budget of iterations.
3.4 Increasing Plan and Code Diversity
We use the following two techniques to boost the diversity in plan (and in the case of LLM ensembling, code) generation and prevent the model from repeatedly selecting the same optimization:
Optimization Menu Dropout. Inspired by methods for preventing neural networks from overfitting srivastava2014dropout, we implement dropout in our optimization menu. Each time a plan is generated, each menu option in the list of optimizations has a chance to be removed.
LLM Ensembling. Ensembling LLMs is known to improve diversity and quality of results jiang2023llmblender. To further increase the diversity of generated plans and code, whenever multiple candidates are sampled, we divide these requests between different LLMs.
We ablate these techniques, along with other components of Autocomp, in Appendix B.
3.5 Schedule Reuse
Running Autocomp’s search starting with unoptimized code results in optimized code that outperforms all prior approaches, as we will discuss in Sec.˜4. However, using this approach with every new software workload can be costly, as Autocomp involves multiple LLM invocations and hardware simulations. A natural question, then, is whether the schedules discovered for one workload can be used to accelerate the optimization of others. We draw inspiration from traditional optimized libraries like BLAS blackford2002blas, where hand-tuned schedules are reused across GEMM shapes, and extend Autocomp with schedule reuse.
To do so, we first record the best known schedule for a particular tensor operation. Then, during planning for new GEMMs with the same aspect ratios or with two shared dimensions, rather than exploring the full menu, we prompt the LLM to apply the menu options used in our recorded schedule, one at a time. As we are not exploring the full menu, we can use a smaller beam width and sample count, reducing both LLM calls and search time. After completing this lightweight search, we take the best-performing code so far and further refine it by invoking the full Autocomp search for a small number of iterations. This resembles the classic exploration-exploitation trade-off in optimization: by reusing a schedule we exploit a known high-quality schedule and avoid the initial exploration cost for a new workload.
4 Evaluation
4.1 Hardware Platforms
4.1.1 Gemmini
Gemmini gemmini is an accelerator generator that can generate systolic array- and vector-style tensor accelerators with a wide range of data types and sizes. Gemmini is ideal for evaluating Autocomp as it: 1) generates accelerators that deliver performance comparable to commercial ones, 2) is open-source, enabling instantiation of different accelerator instances, user modifications to the software toolchain, and extraction of fine-grained performance feedback, and 3) supports fast and cycle-accurate hardware simulation via FireSim firesim. The specialized and low-resource nature of generated Gemmini instances is ideal for our evaluation, as it eliminates data contamination and makes directly prompting LLMs challenging.
4.1.2 AWS Trainium
Trainium trainium is a family of state-of-the-art tensor accelerators built and deployed by Amazon Web Services (AWS). Trainium’s software stack includes several different entry points for users, including PyTorch, JAX, and the Neuron Kernel Interface (NKI). In our evaluation, we optimize code written in NKI, which enables lower-level control of computation and data movement. While Trainium accelerators are a real-world, high-performance industry backend, they are very low-resource, as Trainium was first deployed in 2022 trainium_released with NKI only being released in 2024 trainium_nki_released. This makes Trainium an ideal target for evaluating Autocomp. We optimize code for Trainium 1 (specifically, a trn1.2xlarge instance) as later generations are not yet widely available. This instance contains two NeuronCore-v2, each of which contains scalar, vector, and tensor (systolic array) engines, as well as on-chip scratchpad and accumulator memories (called SBUF and PSUM), which communicate with main memory, and supports a wide range of data types.
4.1.3 NVIDIA L40S GPU
As discussed in Sec.˜1, ML-based compilers such as TVM Ansor ansor have been developed to generate high-performance deep learning code for CPUs and GPUs. However, these compilers do not support low-resource accelerators such as Gemmini and Trainium, and porting TVM to these platforms would require huge amounts of engineering effort. As a result, we additionally evaluate Autocomp on a GPU backend (specifically, NVIDIA L40S) to compare Autocomp’s performance against TVM’s ML-based auto-scheduler and further demonstrate Autocomp’s portability and effectiveness. The L40S is a modern datacenter GPU that includes Tensor Cores specialized for tensor processing, and supports a wide range of data types and CUDA libraries.
4.2 Discussion on Autocomp’s Portability
Autocomp’s initial implementation targeted Gemmini. Once this initial implementation was complete, it was remarkably easy to port Autocomp to new hardware platforms. Qualitatively, Autocomp is easily portable thanks to the fact that for new hardware platforms, only prompts need to be changed—specifically the optimization menu and ISA description need to be updated—there is no complex compiler backend code to write. Quantitatively, we were able to get up and running with each of the Trainium and GPU backends in less than a day’s effort by one graduate student. When compared against the large industrial efforts behind multi-platform compilers such as XLA, this miniscule investment clearly illustrates Autocomp’s extreme portability and ease of use.
| Baseline |
Gemmini (4x4 FP32) |
Gemmini SW lib | ✗ | PyTorch NeuronX trainium_xla | PyTorch | |||
| Unoptimized low-level | Exo unoptimized | dong2024designspaceexplorationembedded |
| Optimized low-level |
| ML cost model | ✗ | ✗ | ✗ |
| Hardware FSM |
4.3 Baselines
We compare Autocomp with the following baselines. Note that not all baselines are available for each hardware platform.
1.
High-Level Software Library. Gemmini ships with a software library that uses heuristics to tile and run GEMMs and convolutions on generated accelerators. As loop ordering is fixed and loop bounds, addresses, and indices must be computed at runtime, this implementation incurs significant software overhead and cannot fully utilize hardware resources. Trainium can be run using PyTorch as a frontend, but similarly, this prevents the user from implementing low-level optimizations. Instead, Trainium’s NeuronX compiler, based on the XLA compiler xla; trainium_xla, can automatically optimize a PyTorch module by tracing it and taking advantage of fixed shapes and fixed control flows to produce a fused computation graph. Finally, NVIDIA GPUs are commonly used with PyTorch as a frontend for high-level programming.
2.
Unoptimized Low-Level Code. Exo exo is a DSL for tensor computation and scheduling. It comes with a basic compiler that emits GEMM or convolution code in Gemmini’s ISA. However, without benchmark-specific optimization, performance is highly suboptimal, as hardware resources such as local memories tend to be underutilized. Trainium provides the nki-samples repository, which contains naive, unoptimized NKI implementations of several tensor operations in the directory src/nki_samples/tutorial.
3.
Hand-Optimized Low-Level Code. In exo, Exo’s and Gemmini’s developers spent significant effort manually writing and hand-tuning benchmark- and accelerator-specific schedules for each of the GEMM and convolution sizes in Sec.˜4.3 and 4.3. This is the previous best known software implementation for these benchmarks. Similarly, Trainium provides optimized versions of the naive tutorial implementations in nki-samples. In addition, nki-samples provides a set of advanced implementations that “showcase cutting-edge optimizations and specialized implementations,” in the directory contributed/neuron-team-kernels.
4.
Machine Learning-Based Cost Model. We use TVM MetaSchedule as our baseline for ML-based deep learning compilers for GPUs. MetaSchedule is TVM’s latest iteration of auto-scheduler, succeeding TVM Ansor. This auto-scheduler runs a large number of samples, selected via evolutionary search, to train a cost model based on a large number of hand-crafted features shao2022metaschedule. No such ML-based compiler exists for the low-resource accelerators, Gemmini and Trainium.
5.
Gemmini Hardware FSM. Gemmini can generate accelerators with specialized hardware units for two coarse-grained operations: GEMM and convolution. These hardware units, implemented as finite state machines (FSMs), encode control sequences for each of these operations in hardware. If tuned correctly, the FSMs can exceed the theoretical maximum performance of any software-based implementation, as hardware is inherently parallel. However, this is accomplished at the cost of scheduling flexibility, as well as increased area, power, and hardware complexity. We use the hardware FSM as a reference for the highest achievable performance, but do not expect to exceed its performance for these GEMM/convolution benchmarks, as its compute utilization exceeds 90% for all but one benchmark.
4.4 Gemmini Evaluation
On Gemmini-generated accelerators, we evaluate the effectiveness of Autocomp on three distinct types of workloads: 1) matrix multiplication (GEMM) derived from ResNet-50, 2) convolution derived from ResNet-50, and 3) robotics code used for model-predictive control. We ensemble OpenAI’s o3-mini and gpt-4o (via the OpenAI API Platform) for both phases, with temperature 1.0. Menu options are dropped out with 70% probability. During the Implement phase of prompting, in cases where the optimization plan contains the string ‘‘tiling’’, we provide an ICL Example example of code (from a different workload) before and after changing one tiling factor, as tiling is a key optimization that requires modifications across the program, making it challenging to implement.
For GEMM and convolution, we use unoptimized low-level code generated by Exo, which contains statically computed loops, addresses, and indices, as Autocomp’s starting point. This simplifies code generation and allows us to directly compare the effectiveness of Autocomp to hand-optimization. The target for these workloads is a Gemmini-generated accelerator with 1616 systolic array, 8-bit integer data type (accumulating in 32-bit), 256 KB scratchpad, and 64 KB accumulator, as used by exo.
The third workload type, fine-grained linear algebra, contains sequences of element-wise operations and matrix-vector multiplications from robotics control. As this is not directly supported by Exo, we compare to an unoptimized software implementation ported to accelerator code by dong2024designspaceexplorationembedded (also used as the starting point for Autocomp), and an extensively hand-optimized hardware FSM-based implementation written by an expert. These are run on a 32-bit floating point accelerator with a 44 systolic array, and the same scratchpad and accumulator sizes as above, as used by dong2024designspaceexplorationembedded.
4.4.1 Matrix multiplication (GEMM)
We run Autocomp on a set of GEMM benchmarks selected by exo from ResNet-50 resnet for diversity in size and shape. We run search with beam size , plans per element in the beam, code candidates per plan, and iterations.
Sec.˜4.3 shows that Autocomp significantly outperforms even extensively hand-optimized code (Exo Opt) by a geomean of 1.4, Exo Unoptimized code (the starting point of Autocomp’s search) by 2.9, and Gemmini’s software library by 5.6. Autocomp consistently achieves at least 85% of the hardware FSM’s utilization (91% on average).
Autocomp especially outperforms prior implementations thanks to extensive exploration of software pipelining and double-buffering, which allows better overlapping of data movement and computation, for example by double-buffering both the scratchpad and accumulator. In many cases, Autocomp’s exploration also leads to different tiling and loop ordering choices than hand-optimized code, reducing data movement. We analyze Autocomp-generated GEMM code in further detail in Appendix D.
4.4.2 Convolution
We also optimize convolution benchmarks from ResNet-50 via the same process. Compared to the GEMM benchmarks, this code contains more loops and is more complex. In this case, we run beam search with beam size , plans, and code candidates, for iterations.
Compared to GEMM, convolution provides less room for improvement over both the Gemmini software library and exo’s implementation, as even the hardware FSM only achieves a 3.3 geomean speedup over the software library, compared to 6.1 for GEMM. This is because on average the Gemmini software library achieves 28% of the theoretical maximum compute utilization, compared to 16% for GEMM. As discussed by gemmini, at ResNet-50’s tensor sizes, convolutions have greater arithmetic intensity than GEMMs, making them less memory-bound and causing the Gemmini software library’s suboptimal data orchestration to be less impactful.
Nonetheless, as shown in Sec.˜4.3, Autocomp still exceeds the previous best known hand-optimized software ISA-based implementation (Exo Opt) by up to 1.2 and by a geomean of 1.1, via similar strategies as for GEMM. It also outperforms Exo Unoptimized code by 2.9, and Gemmini’s software library by 2.6, and in all cases achieves at least 78% of the hardware FSM’s utilization.
4.4.3 Fine-Grained Linear Algebra
Finally, we optimize fine-grained linear algebra benchmarks from the TinyMPC model-predictive control library nguyen2024tinympc, specifically the forward and backward passes of the primal update step. These benchmarks contain sequences of floating-point matrix-vector multiplications, interleaved with element-wise addition and subtraction. The inclusion of CPU-accelerator dependencies, low reuse, and a high ratio of data movement to computation leads to low accelerator utilization and makes this code challenging to optimize.
We compare Autocomp-generated code against dong2024designspaceexplorationembedded’s unoptimized software-based implementation on a 44 FP32 accelerator. For this work, we additionally had an expert hand-tune a hardware FSM-based implementation. The unoptimized software-based implementation is used as the starting point for search, and we use the same search parameters as for convolution, except with iterations. Some of the optimization menu options are different from those used for GEMM/convolution (see Appendix C). As shown in Fig.˜9, Autocomp outperforms even the expert-optimized hardware FSM implementation on the forward pass (by 1.6), and across benchmarks speeds up unoptimized code by a geomean of 2.7.
To outperform the hardware FSM implementation, Autocomp harnesses the flexibility of software-based implementation. It optimizes the code by hoisting data loads shared between kernels (reducing data movement beyond what is possible for the hardware FSM implementation), as well as utilizing fine-grained software pipelining and eliminating blocking operations where possible. This experiment highlights Autocomp’s adaptability: we optimize a new benchmark, running on an accelerator with a new size and data type, with highly different performance characteristics from previous experiments, by changing a few lines across the Accelerator ISA and Optimization Menu sections of the prompt.
4.5 Trainium Evaluation
As discussed above, we optimize two categories of Trainium benchmarks: Tutorial (starting from naive code and comparing to optimized code) and Advanced (starting from optimized code). Because Trainium has a larger ISA than Gemmini, for each benchmark, we manually specify relevant instructions, and include a description and examples (sourced from Trainium’s documentation222https://awsdocs-neuron.readthedocs-hosted.com/en/latest/nki/index.html) for only those instructions in the Accelerator ISA section of the prompt. Furthermore, as Trainium’s compiler provides relatively good error messages for syntax errors, when a code implementation fails correctness checking, we prompt the LLM with the Accelerator ISA, the code, and the compiler-generated error message, and then evaluate the fixed code again.
For all experiments, we ensemble OpenAI’s o4-mini and gpt-5 for both phases. We search with 70% menu dropout, beam size , plans, code candidates, and iterations.
4.5.1 Tutorial Workloads
As discussed in Sec.˜4.3, AWS provides a set of tutorial NKI implementations that demonstrate varying levels of optimization. These workloads include key deep learning operators of varying scopes, detailed in Appendix A. For these workloads, we start optimization from the unoptimized naive NKI implementation, if one is available (for RMSNorm and Stable Diffusion attention, we start from the optimized implementation).
Many of the optimizations in our Optimization Menu for Trainium are based on these tutorials, so we expect Autocomp to at least match the performance of the fully optimized code. Autocomp not only does so, but as shown in Fig.˜10, outperforms hand-optimized code by a geomean of 1.36. In doing so, Autocomp speeds up the starting code used as input (either unoptimized or optimized NKI code) by a geomean of 2.51.
The nki-samples repository also contains PyTorch implementations of these operators, which we compile with Trainium’s NeuronX compiler. Autocomp-generated code outperforms the code compiled from PyTorch by 13.52.
4.5.2 Advanced Workloads
AWS also provides a set of highly optimized NKI implementations written by expert kernel engineers. For these workloads, we start search from the already optimized code. Since these workloads are already optimized, any improvement is a highly positive result. As shown in Fig.˜11, we find that Autocomp is able to optimize these workloads by a geomean of 1.9, including speeding up 2D convolution by 2.98 and causal self-attention by up to 1.36. Unfortunately, no matching PyTorch implementation is provided for these workloads.
Notably, Autocomp speeds up 1D depthwise convolution by 17.37. It does so through a sequence of optimizations that takes advantage of both the specific target shape as well as the code’s inherent inefficiencies: first, it decreases allocated tile sizes in the scratchpad to prevent spilling to main memory, which allows the next iteration to move the new smaller tile-sized accumulations into the accumulator. Then, it swaps loop ordering to increase filter reuse over batches, and finally, re-expands scratchpad tile size by adding a new level of tiling over the long output dimension, increasing data reuse in the scratchpad.
4.6 NVIDIA L40S GPU Evaluation
When optimizing GPU code, we remove the Accelerator ISA component of the prompt, as Python/CUDA are relatively high-resource languages compared to Gemmini and Trainium. The Optimization Menu is crafted from a variety of sources mills_2024; lange2025robustagenticcudakernel; li2025cudal1improvingcudaoptimization; ouyang_liang_mirhoseini_2025. For example, “Use shared memory to reduce global memory bandwidth usage,” or “Minimize divergent branches within warps” (the full list can be found in Appendix C). We increase menu option dropout to 80% due to the large resulting menu. We ensemble OpenAI’s o4-mini and gpt-5 for both phases, with temperature 1.0. We use beam size , plans, code candidates, and iterations. During the Implement phase, when the string ‘‘tensor core’’ is detected in a plan, we include several ICL Examples of tensor core usage in CUDA, sourced from CUDA/cuBLAS/cuDNN documentation.
For our evaluation, we select a few kernels from KernelBench ouyang2025kernelbench. We focus on kernels from level 1 (individual operators) which do tensor-tensor computation, as we found that with more complex operations, porting PyTorch code to TVM was highly cumbersome and TVM’s scheduler failed at a high rate or was extremely slow. We tune TVM MetaSchedule with 1000 trials.
KernelBench uses PyTorch code as a starting point for optimization. As a result, in the first two iterations of search, we focus on generating an initial inline CUDA-based implementation using a reduced menu (see Appendix C).
As shown in Fig.˜12, Autocomp generates code that is faster than PyTorch on every benchmark, speeding up the benchmarks by a geomean of 2.05, whereas TVM only outperforms PyTorch on two benchmarks, and on average is 0.54 as fast as PyTorch. TVM narrowly outperforms Autocomp on only benchmark 9, a GEMM with a small reduction dimension. One key to Autocomp’s performance on these benchmarks is its ability to explore different methods of enabling Tensor Core utilization, such as through PyTorch APIs and CUDA libraries such as cuBLASLt, and identify the best-performing implementation that passes KernelBench’s accuracy check. On the other hand, since KernelBench focuses on 32-bit floating point operators, TVM MetaSchedule is unable to generate Tensor Core code without manual intervention.
These experiments show that not only is Autocomp more portable across hardware platforms and easier to use than TVM’s auto-scheduler (since it can directly optimize PyTorch code), it is also significantly more performant for tensor operations from KernelBench’s level 1.
5 Improving Sample Efficiency With Schedule Reuse
In this experiment, we illustrate that optimization schedules generated by Autocomp for one tensor operation can be effectively reused to accelerate the optimization of similar tensor operations (as described in Sec.˜3.5). For this experiment, we generate additional benchmarks structurally similar (either same aspect ratio or two shared dimensions) to those in Sec.˜4.4.1. For each new benchmark, in the plan reuse phase, we search with beam size , plans, and code candidates. Once the reuse sequence is exhausted, we switch to a full Autocomp search with beam size , plans, and code candidates.
We consider GEMMs from the following categories: 1) square, 2) column-dominant and, 3) row-dominant, with dimensions specified in Table˜2. The “base” benchmark is optimized using full search, i.e., without any schedule reuse, and its Autocomp generated schedule is then reused to optimize the other two benchmarks in its category.
| Category | Base Benchmark | Reuse Targets |
| Square | 1024×1024×1024 | 512×512×512, 256×256×256 |
| Column-dominant | 12544×256x64 | 6272×256x64, 12544×128x64 |
| Row-dominant | 128×1024×1024 | 64×1024×1024, 32×1024×1024 |
We include the full optimization traces for Autocomp with and without reuse for one of these benchmarks: 512×512×512 (square GEMM), reused from a 1024x1024x1024 GEMM. As shown in Fig.˜13, with reuse, Autocomp consistently achieves better speedups with the same number of samples.
For the remaining benchmarks, we perform an iso-samples comparison: we fix the total number of LLM calls and compare reuse to full search. In Fig.˜13, the red vertical lines compare the iso-sample performance for our two examples. Across all reuse targets in Table˜2, with a budget of 100 samples (10% of the calls used by full search), Autocomp achieves geomean speedups over Gemmini’s software library of with reuse, compared to without reuse. With a 200 call budget, we observe speedup with reuse, compared to without reuse. This demonstrates Autocomp schedules are generalizable and can help reduce search cost when optimizing similar benchmarks.
6 Conclusion
In this paper, we demonstrate how to construct an LLM-based flow to automatically optimize low-resource accelerator code at superhuman levels. Across three distinct hardware platforms, Autocomp generates code that significantly outperforms baselines on real-world benchmarks, while requiring far less manual effort than implementing a traditional compiler or hand-optimizing code. This demonstrates the effectiveness of Autocomp’s approach and its potential to serve as a key component in the accelerator design process.
Appendix A Benchmark Details
A.1 Gemmini Experiments
We have uploaded the Gemmini benchmarks to Google Drive333https://drive.google.com/drive/folders/1eceVTv7noM6Rp63ap9PTL5t6qMYOlfcr?. Gemmini’s software library is from the gemmini-rocc-tests repository444https://github.com/ucb-bar/gemmini-rocc-tests/tree/1a1a1c6bd60df6d7cae3d87aac96c8f406cae084. Exo optimized code was generated using using the published schedules in Exo’s GitHub repository555https://github.com/exo-lang/exo/tree/4c65f576b296025b8dcd86b2c4e41769507969cf/apps/gemmini/src/exo.
A.2 Trainium Experiments
| Operator | Configuration |
| RMSNorm | |
| LayerNorm | |
| GEMM | |
| Mamba | , , , |
| Self-Attention | , |
| Stable Diffusion Attention | , |
| Operator | Configuration |
| Cumsum | |
| Transpose | |
| Max Pooling | , pool |
| RoPE | |
| Depthwise Conv1D | (NCHW) |
| Conv2D | batch size , input , kernel , input channels , output channels |
| Causal Self-Attention seq=2048 | , |
| Causal Self-Attention seq=16384 | , , |
| Multi-head Causal Self-Attention | , , |
Tables˜3 and 4 detail the shapes and configurations used for our Trainium evaluation. Reference code was directly copied from the nki-samples repository666https://github.com/aws-neuron/nki-samples/tree/5cfeabca92a64f642a154ef835cbcede609016a3, with some implementations requiring small modifications to run in isolation on a trn1.2xlarge instance. We attempted to use the same shapes as in the original nki-samples code, whenever examples were provided and the shapes were large enough to show meaningful performance improvements. For causal self-attention with seq=16384, we measure the latency of calculating self attention between a query tile of length with key/value arrays of length .
A.3 NVIDIA L40S GPU Experiments
| Benchmark # | Configuration |
| 1 | square GEMM |
| 2 | GEMM |
| 6 | large K GEMM |
| 9 | small K GEMM |
| 10 | 3DMM |
| 50 | Conv2D weight , input , output (stride 4, padding 2) |
| 54 | Conv3D weight , input , output (stride 1, padding 0) |
| 56 | Conv2D asymmetric weight , asymmetric input , output (stride 1, padding 0) |
Table˜5 lists the details of the KernelBench benchmarks selected for our GPU evaluation. We use shapes from KernelBench v0.1. The exact PyTorch operators used can be found in the KernelBench repository777https://github.com/ScalingIntelligence/KernelBench/tree/6b1cea28ddea79af6b295fb9a3e3413f393efe3d/KernelBench/level1.
Appendix B Ablation Studies
| Experiment |
| Baseline (Exo Unopt) | 1.67 | 0.91 | ||
| No Accelerator ISA | 3.11 | 2.51 | ||
| No Optimization Menu | 2.34 | 0.97 | ||
| No Optimization Menu Dropout | 4.72 | 2.30 | ||
| No LLM Ensemble (o3-mini only) | 4.67 | 2.08 | ||
| No Hardware Perf Feedback | 4.91 | 2.61 | ||
| LLM Selection (DeepSeek-R1) | 4.89 | 2.25 | ||
| LLM Selection (Gemini 2.5 Flash) | 3.63 | 2.42 | ||
| LLM Selection (Llama 4 Maverick) | 4.14 | 1.01 | ||
| Autocomp | 5.53 | 2.72 | ||
| Variance Checking | 5.20.33 | 2.610.11 |
In this section, we ablate various features of Autocomp to investigate their effect on optimization performance. We focus on two specific benchmarks from Sec.˜4—our 12544x64x256 GEMM and our 4x3x14x256x256 convolution on a 16x16 INT8 Gemmini—to isolate the effects of these ablations while limiting the cost of running this extensive exploration.
B.1 Accelerator ISA
We find that for GEMM and convolution code, removing the ISA significantly deteriorates performance. Still, Autocomp is able to improve over the original code by a notable margin even without the ISA, given that all its other features are still enabled (see Table˜6). This is because we inherently provide an example of accelerator ISA code at each step via the current code, so the model is able to infer some properties of the accelerator ISA. In addition, many of the nested loop optimizations for the GEMM and convolution workloads are well-understood transformations that operate on the C-syntax loops, addresses, and indices in the code, which resemble general-purpose programming, rather than using accelerator-specific constructs such as configuration instructions. However, full Autocomp performance is not matched as the proportion of functionally correct responses is lower, and instruction-level optimizations cannot easily be identified. For example, the first-compute handling optimization in Appendix D.1’s GEMM example and the negative-scaled bias loading in Appendix D.2’s fine-grained linear algebra example would not have been identified without knowledge of the ISA. Overall, we find that the accelerator ISA is an important part of Autocomp’s prompts.
B.2 Optimization Menu
We ablate the optimization menu by removing the menu in the planning prompt, and instead simply asking the model to select one optimization and generate a plan. From this experiment, we find that adding domain knowledge and optimization diversity via the optimization menu is essential to Autocomp. As shown in Table˜6, optimization performance completely deteriorates without the optimization menu. Qualitatively, without the optimization menu, we find that the models tend to repeat similar optimizations, with significantly less diversity and relevance in the generated optimization plans.
B.3 Optimization Menu Dropout
“Dropout” for optimization menu options is a key contribution of this work that increases the diversity of generated optimization plans. Table˜6 shows that menu dropout has a significant effect on performance. Qualitatively, we find that without dropout, models tend to be biased towards a limited set of menu options, a limitation which can be resolved via menu dropout.
B.4 Hardware Performance Feedback
As discussed in Sec.˜3.2, during plan generation, we include the latency, scratchpad utilization, and accumulator utilization of the original code. Table˜6 shows that this component is helpful, but in some cases its effects may be limited. This is because the options listed in the optimization menu already capture some of the metrics measured in our performance feedback, for example the menu options which suggest using larger tile sizes. Hardware feedback such as scratchpad and accumulator utilization only serves to augment elaboration of these menu options by providing exact measurements.
B.5 LLM Ensembling
Splitting requests between LLMs in an ensemble also encourages diversity of generated plans and code. Qualitatively, we find that the responses, especially during the planning phase, generated by different LLMs differ substantially. Our experiments in Table˜6 show that using individual models, such o3-mini or DeepSeek-R1, on their own results in significantly lower performance.
B.6 LLM Selection
In Sec.˜4, we use an ensemble of gpt-4o and o3-mini for our search. To demonstrate that Autocomp does not depend on a particular family of models, we run Autocomp with several alternative LLMs on the same benchmarks used for other ablation experiments above. In addition to DeepSeek-R1 deepseekr1, we run Autocomp with two smaller models, Gemini 2.5 Flash and Llama 4 Maverick. For these experiments, each model was used for both the planning and code generation phases without ensembling, keeping the search parameters identical to those used for matrix multiplication and convolution in Sec.˜4. As shown in Table˜6, Autocomp with DeepSeek-R1 is able to optimize both GEMM and convolution, achieving substantial speed-ups over the unoptimized code. Even with smaller models, Autocomp shows substantial speed-ups. Similarly to the LLM ensembling ablation study above, gains are slightly smaller than when all techniques are applied. Nonetheless, this demonstrates that Autocomp is efficient and flexible across different LLMs.
B.7 Variance Checking
To account for the stochasticity in Autocomp’s search, we ran Autocomp on three different seeds using the same benchmarks as those used for other ablation experiments. In Table˜6, we report the mean and standard deviation of the final speedups. The performance spread was minimal, with the lowest-performing run for GEMM still achieving a speedup within 11.8% of the peak result, and within 8.5% for convolution. This demonstrates that Autocomp’s search is robust and consistently finds high-performing solutions.
Appendix C Prompts
Appendix D Code Examples
In this section, we discuss in greater depth what optimizations Autocomp applies in our evaluations and how Autocomp is able to achieve significantly better performance than hand-optimized code.
D.1 16x16 INT8 Gemmini—12544x64x256 GEMM
Fig.˜28 contains the unoptimized Exo-generated code, used as the starting point for search. Fig.˜29 contains the code generated by Exo after hand-optimization by exo. Figs.˜30, 31 and 32 contain the result of Autocomp optimization on Exo Unoptimized code. While the code is substantially transformed from the original code, some aspects remain the same. For example, in this case the configuration instructions and loop ordering remain largely the same.
Of course, many optimizations have been applied to the code. We briefly summarize the optimization menu options selected and plans generated during the optimization process for this code. We also include the speedup after each respective optimization.
1.
1.67: initial speedup of Exo Unoptimized code over Gemmini’s software library before any optimization.
2.
1.93 after “hoist redundant operations out of loops”. This plan hoists constants like tile_dim = 16 and loop-invariant expressions like ko * 64 and k * 16 out of inner loops. These precomputed values are reused inside Gemmini ops in each iterations, so we should reduce the number of times they must be calculated.
3.
1.95 after “double buffering”. This plan defines two buffer regions for matrices A and B in the scratchpad. A buffer_toggle flag is introduced to alternate between these buffers each iteration. All mvin, preload, and compute instructions are updated to use the active buffer based on the toggle. Data loading for the next iteration is scheduled earlier to overlap with current computation. Address calculations are adjusted to include buffer offsets accordingly.
4.
2.15 after “pipeline operations to better overlap computation and data movement”. Moves mvin2 (A tile load) to immediately after compute_preloaded in the ko loop to overlap A prefetch with current compute. Moves mvin3 (B tile load) earlier in the k loop, before the next compute, to overlap B prefetch with current compute.
5.
3.13 after “load data to the scratchpad across outer loop iterations and use if statements to prevent redundant loads on loops inner to those”. Adds a one-time load of the entire B matrix into a new scratchpad region (specifically, new_B_base = 8192) before the main i loop. Replaces repeated mvin3 calls in inner loops with offset calculations into this preloaded B region.
6.
3.54 after “move more data to the scratchpad in a more outer loop to increase data reuse”. Before the reduction (k) loop, load the entire 16x256 A-tile into scratchpad using four mvin2 calls to a new A_tile_base region. Remove the double-buffering (a_toggle) for A since the full tile is now resident. Replace nested ko and k loops with a single loop over 16 segments, computing each 16x16 A sub-tile address from A_tile_base. Adjust B tile preload and compute_preloaded calls accordingly, using a simpler tile index.
7.
4.87 after “double buffering”. Restores double buffering by reserving two scratchpad regions: A_tile_base0 and A_tile_base1, each holding one full 16x256 A-tile. In the outer i loop, alternate buffers using i % 2 to select current_buffer and next_buffer. In each iteration (except last), while computing with current_buffer, issue mvin2 to load the next A-tile into next_buffer. In the inner compute loops, use current_buffer for A-tile addresses. After compute and mvout, the next iteration uses the preloaded data.
8.
5.21 after “double buffering”. Double-buffers accumulator by allocating two accumulator regions: acc_base0 and acc_base1. In each i iteration, compute into cur_acc_base and mvout from prev_acc_base (except on first iteration), and swap cur_acc_base and prev_acc_base at the end of the loop.
9.
5.23 after “loop unrolling”. Unrolls the innermost loop (j_in_o) by a factor of 4.
10.
5.53 after “fuse loops”. “Fuses” loops by eliminating the loop over j_in_o where we mvin 0s to the accumulator. Instead, use the ability of preload to overwrite the values in the accumulator rather than accumulating, when beginning a new partial sum.
From this example, we observe that a diverse set of optimizations is selected, and that speedups are distributed throughout the optimization process rather than concentrated in just one or two steps, showing the importance of a well-designed iterative search process. From here, we summarize the differences between Autocomp-generated code and the previous best code (Exo Opt):
Tiling. The Exo Opt code loads 128256 tiles of A, whereas the Autocomp-generated code loads 32256 tiles (divided into two 16256 tiles) of A. While this means there is less reuse for the Autocomp-generated code, there is also less overhead needed for compute instructions to wait for each A tile to be loaded to the scratchpad. In combination with the rest of the optimizations applied by Autocomp, this leads to improved performance.
Double-buffering. In the Autocomp-generated code, we see that both the scratchpad and accumulator are explicitly double-buffered. In the schedule, we can see that double buffering is applied 3 times. Initially (in step 3), both the A and B matrices (where matrix multiplication is represented as ABC), are double buffered in the scratchpad. However, after steps 5 and 6, B and A (respectively) are no longer double-buffered as larger tiles are loaded before beginning computation. The accumulator is double-buffered in step 8, resulting in the code below. The Exo Opt code relies on the accelerator’s out-of-order execution to handle executing mvin and mvout instructions without dependencies, ahead of the order in which they are issued.
Software pipelining. The Autocomp-generated code explicitly issues A mvin instructions before they are needed for computation, whereas as above the Exo Opt code relies on hardware to handle overlapping of data movement and compute. Also, the Autocomp-generated code explictly issues all B mvin instructions at the beginning of the program, whereas the Exo Opt code interleaves these instructions with computation (but still loads the entire B matrix to the scratchpad, once overall). This does not have a significant impact on performance, but LLM-generated code is naturally biased towards such an implementation due to its simplicity.
First-compute handling. The Autocomp-generated code utilizes the ability of compute instructions to overwrite the accumulator, whereas Exo Opt code explicitly issues mvin instructions to zero out the accumulator before beginning computation on a tile.
Arithmetic simplification. Arithmetic on constants is fully simplified and handled inside shared variables wherever possible in the Autocomp-generated code, reducing the overhead of non-accelerator instructions.
Overall, we find that compared to the Exo Opt code, Autocomp-generated code applies more techniques to minimize the amount of CPU overhead during execution. The smaller tiles it uses, in combination with its more explicit application of double-buffering and software pipelining, results in highly tuned, fine-grained overlapping of data movement and computation and a very high level of performance.
D.2 4x4 FP32 Gemmini—TinyMPC Primal Update Forward Pass (Fine-Grained Linear Algebra)
The forward pass of the TinyMPC primal update step computes the following operations:
Where is a 12x12 matrix, is a 12x4 matrix, is a 4x12 matrix, is an NHORIZONx12 matrix (where individual columns are accessed here via indexing), is an NHORIZONx4 matrix, and is an NHORIZONx4 matrix. , , , and the 0th column of are inputs, and is the output. The 1st to (NHORIZON-1)th column of are intermediate values computed over the course of the benchmark.
This process is repeated until a time horizon, defined as NHORIZON in our code and set to 5 for our evaluations. Note that is allocated as an (NHORIZON+1)x12 matrix in our code since the unoptimized code accesses up to the (NHORIZON)th column.
Autocomp generates the code in Fig.˜36, optimized over several steps from the starting code in Fig.˜28. The following optimizations are applied, with the following speedups after each optimization:
1.
1: in this case we treat the unoptimized software as the baseline for speedup, so by definition its speedup is 1.
2.
1.07 after “hoist redundant operations out of loops”. Hoists the mvin calls for the constant matrices Kinf, Adyn, and Bdyn above the NHORIZON loop and executes them once rather than in every iteration. Any associated config_ex and config_ld calls are also moved outside the loop if needed. The compute calls use the same scratchpad addresses and may set B_spad_addr = 0xffffffff to indicate that the weights are already loaded.
3.
1.13 after “loop reordering”. This plan claims to move the configuration and mvin instructions for Kinf, Adyn, and Bdyn before the NHORIZON loop, but this has already been handled in the previous step. In reality, only some configuration instructions that were unnecessarily left behind by the previous step are hoisted.
4.
2.00 after “move CPU-based computation to the accelerator”. Replaces CPU-based element-wise negation and addition with equivalent Gemmini compute instructions. Specifically, when x_i and d_i are loaded, a scaling factor of -1 is used to negate it in the scratchpad. The product of Kinf and x_i is kept in the accumulator and the negated d_i is multiplied by a dummy 0 vector in order accumulate it on top of this product, enabling data to be kept resident in the accumulator rather than moving it back and forth between the CPU and accelerator memory.
5.
2.12 after “hoisting redundant operations out of loops”. Identifies a few config_ld instructions that are still redundant and removes them from the loop.
6.
2.66 after “loop unrolling”. Changes the outer loop to increment by 2 instead of 1 and duplicate the loop body so each iteration computes two time steps: first u[i] and x[i+1], then u[i+1] and x[i+2]. Also merges fence calls at the end of the unrolled loop body if possible. The implementation very aggressively removes fence instructions, which is actually correct as reuse of the same accelerator memory instructions by subsequent mvin and mvout instructions means that dependencies can be handled internally to the accelerator, rather than via synchronization of both the CPU and accelerator.
7.
2.95 after “loop fusion”. Undoes the loop unrolling from the previous step, but keeps the reduced fence instructions. Additionally, this eliminates the unnecessary calculation of x[NHORIZON] during execution, which saves cycles. This optimization makes sense since it is likely the reduced number of fence instructions that improved performance in step 6, rather than the loop unrolling (which usually provides limited benefit, if any).
We further discuss the similarities and differences between the hand-optimized hardware FSM-based implementation in Fig.˜35 and the Autocomp-optimized code:
Data orchestration. Due to the coarse-grained nature of the hardware FSM operations and the fact that data movement between the accelerator and main memory is handled within hardware, we are not able to hoist shared data loads of the Kinf, Adyn, and Bdyn matrices out of the loop. This is the main advantage the Autocomp-generated software ISA-based implementation has over the hardware FSM-based implementation.
Operator fusion. Both implementations are able to handle all computation on the accelerator, but notably the hardware FSM-based implementation fuses the addition of d[i] into the bias when multiplying Kinf and x[i] (then the whole result is negated while storing to main memory), whereas the addition of d[i] is handled as a separate compute instruction in the Autocomp-generated code. So, the hardware FSM-based implementation actually has an advantage in this regard and the Autocomp-generated code has further room for improvement. Both implementations make use of negative scaling of loaded data (via config_ld instructions) in order to handle subtraction.
Fence instructions. The Autocomp-generated code is able to remove CPU-based fence instructions and instead handle dependencies within the accelerator, whereas the hardware FSM-based code is forced to place a fence after each matrix multiplication, as accumulated results must be moved back to main memory and loaded to the scratchpad for the next operation.
Configuration overhead. While they do not have as much overhead as fence instructions, configuration instructions can also be blocking. The hand-optimized code hoists configuration instructions out of the loop where possible, but since different matrix sizes must be loaded inside the loop, configuration instructions cannot be completely eliminated, giving Autocomp’s code the advantage in this aspect.
Dead code elimination. Both implementations eliminate the extra computation of x[NHORIZON] that is present in the unoptimized code.
Overall, we find that Autocomp identifies all major optimization opportunities available in the code, with the exception of handling subtraction of d[i] slightly suboptimally. Qualitatively, optimizing this code by hand can be difficult due to a lack of readability and the difficulty of debugging low-level accelerator code. The Autocomp-generated code is well-commented and the sequence of optimizations applied can be easily understood. This will be helpful for further optimization of this benchmark or future optimization of other benchmarks, via methods like the schedule reuse demonstrated in Sec.˜5.